# The spectral genus of an isolated hypersurface singularity and a conjecture relating to the Milnor number

### v1

@@ 3. Analytic torsion of $\mathcal{O}_{X}$
$\log\tau(\mathcal{X}_{t},\mathcal{O}_{\mathcal{X}_{t}})=(-1)^{n}\left(\frac{\mu}{(n+2)!}-\widetilde{p}_{g}\right)\log|t|^{2}+o(\log|t|).$
In particular, if the weak form of the conjecture holds, then $\tau(\mathcal{X}_{t},\mathcal{O}_{\mathcal{X}_{t}})^{(-1)^{n}}$ goes to zero as $t$ goes to 0.
Notice that, while the analytic torsion depends on the choice of metrics, the asymptotic behavior is a topological invariant. Also, for families of curves, the $L^{2}$-norm is independent of the choice of auxiliary metric, and the asymptotic behavior is then valid for any choice of Kähler metric on the total space $\mathcal{X}$.
@@
A variant of the above arises from the tautological family $\mathcal{H}$ of hypersurfaces of degree $d$ in $\mathbb{P}^{n+1}$ over the space of parameters $\mathbb{P}^{N}$, where $N={n+1\choose d}-1$. Then $\mathcal{H}$ is a smooth space and inherits a Kähler metric from $\mathbb{P}^{n+1}$ and $\mathbb{P}^{N}$. Hence, the restriction along any curve $\Delta\to\mathbb{P}^{N}$ in which the total space is still smooth is an example of such a degeneration.
@@ 4. Quasi-homogeneous singularities
In this section, we address several cases of the conjecture in the setting of quasi-homogeneous singularities. We also recall some useful notions about Newton polyhedra also utilized in later sections.
@@ 4.1. Newton polyhedra
where the sum runs over integers $k_{i}>0$ such that $\sum k_{i}w_{i}<1$. These are the interior lattice points of the Newton polyhedron of the polynomial $f(x_{0},\ldots,x_{n})$. Eventually, for convenience, we may include the lattice points with $\sum k_{i}w_{i}=1$, since these contribute zero to the sum. For the Milnor number, the formula (4.1) is given in [21, Theorem 1]. The expression (4.2) for $\widetilde{p}_{g}$ can be derived from the computation of the spectral polynomial of $f$ (cf. (2.5) for the definition), due to Steenbrink [32, Example 5.11]:
@@
To this end, it is enough to expand the polynomial into a power series of $T$ with rational exponents, and collect the terms whose exponents are strictly smaller than one. For this purpose, one can ignore the negative $T$ in the numerator, and expand $T^{w_{j}}/(1-T^{w_{j}})=T^{w_{j}}+T^{2w_{j}}+\ldots$. One finds that the exponents $\alpha$ strictly less than one are of the form $\alpha=\sum k_{i}w_{i}$, where the $k_{i}$ run over all the possible integers $k_{i}>0$ such that $\alpha<1$.
@@ 4.3. Homogeneous polynomials
It would be interesting to establish the conjecture for Brieskorn–Pham singularities of the form
@@
corresponding to the choices of weights $w_{i}=1/a_{i}$. To this end, we treat the case when $f$ be a homogenenous polynomial in arbitrary dimension, so that all $w_{i}=\frac{1}{d}$. In this case (4.1) gives
@@
An inductive argument shows that the sum is a polynomial of degree $n+2$ in $d$. Moreover, the sum is empty for $d=0,\ldots,n+1,$ so that the polynomial is of the form $C\cdot d(d-1)\cdots(d-(n+1))$ for some constant $C$. Since the sum is equal to $1$ for $d=n+2$ one sees that $C=\frac{1}{(n+2)!}$. ∎
@@ 4.4. Quasi-homogeneous singularities in dimension one
When $n=1$, the strong form of the conjecture is true in the case of quasi-homogeneous singularities.
In preparation for the proof, we first notice that the invariants $\mu$ and $\widetilde{p}_{g}$ only depend on the analytical type of the germ of the singularity. By [26, Satz 1.3] we can assume that $w_{i}\leq\frac{1}{2}$. Then, by [39], the weights depend only on the topological type of the singularity, which is one of the following:
@@
The case when $a,b$ are relatively prime follows immediately from the computation of Mordell [22, Equation (4)]. The sum in (4.8) is
@@ 5. Irreducible curve singularities
See Némethi [24, Section 3], and in particular Theorem 3.1 therein, and M. Saito [30, Theorem 1.5 & Section 5]. For the spectral genus, we refer to §2.3 above for the expression in terms of Saito’s convention. ∎
@@ 6. Generic singularities with large Newton polyhedra
In this section, we consider isolated hypersurface singularities which are non-degenerate and convenient in the sense of Kouchnirenko [20]. [^2] In op. cit., which is written in French, convenient is called commode. These conditions are formulated in terms of the Newton polyhedron. The non-degeneracy condition is shown to be generic for the Zariski topology. Therefore, we will refer to such singularities as generic. Convenient means that the Newton polyhedron intersects all the coordinate hyperplanes. We refer to op. cit. for details.
@@
Consider now a convenient and generic isolated hypersurface singularity defined by $f\colon(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$. The formula of Kouchnirenko [20, Théorème I] states that
@@
The spectral genus can also be described explicitly in terms of the Newton polyhedron. Let $\Gamma_{c}$ be the compact Newton boundary. It is defined by a homogeneous, concave, piece-wise linear function $\phi$, which takes the value 1 on the faces of $\Gamma_{c}$. The lower Newton polyhedron can then be presented as $\Gamma_{-}=\{x\in\mathbb{R}_{+}^{n+1}\mid\phi(x)\leq 1\}$. Then, by [29], in particular using the formulas on the first page, it follows that the spectral genus is given by
@@
In order to estimate an asymptotic form of the sum (6.2), we will apply a rough version of the Euler–Maclaurin formula for polyhedra. For the statement, we adopt the following convention for Lebesgue measures. Let $P$ be a convex lattice polyhedron in $\mathbb{R}^{d}$, and $Q$ a face of it. Let $\langle Q\rangle_{\mathbb{R}}\subseteq\mathbb{R}^{d}$ be the real vector space parallel to $Q$. Then we denote by $dx$ the Lebesgue measure on $\langle Q\rangle_{\mathbb{R}}$ which gives volume one to the fundamental mesh of the lattice $\mathbb{Z}^{d}\cap\langle Q\rangle_{\mathbb{R}}$.
@@
This is an application of the local Euler–Maclaurin formula of Berline–Vergne [2, Section 5.4, Theorem 5]. The dominant term of the expansion is already given in loc. cit. They discuss the subdominant term in Section 5.5. Notice also that in [2] one supposes that $h$ is a smooth function on $\mathbb{R}^{d}$ with compact support. Since the sum and the first terms of the expansion depend only on the values of $h$ in $P$, which is compact, we can smoothly extend $h$ outside of $P$, still with compact support, for which the formula applies. ∎
@@
In general, the integrals appearing in the previous lemma can be evaluated by applying results of Brion [7, Section 3.2], see also [8, Corollary 6.1.10]. In our setting, it will be enough to have the following.
@@
which concludes the computation. ∎
@@
Let $f(x_{0},\ldots,x_{n})=0$ be a convenient and generic isolated singularity. Then for $k$ sufficiently large, the isolated singularity $f(x_{0}^{k},\ldots,x_{n}^{k})=0$ satisfies the strong form of the conjecture.
@@
where $\phi$ is the piece-wise linear function determining the lower Newton polyhedron $\Gamma_{-}$. To estimate (6.5), we apply Lemma 6.1. For this, we first decompose $\Gamma_{-}$ into simplices with 0 as a vertex and the rest of vertices in $\mathbb{Z}^{n+1}\cap\Gamma_{c}$. These define a simplicial complex, whose top dimensional simplices we denote by $\sigma_{i}$, and whose codimension 1 simplices we denote by $\tau_{j}$. On each simplex, the function $\phi$ is linear. A first application of the lemma gives
@@
For big enough $k$, the leading term, which is positive, dominates and this concludes the proof of the theorem. ∎
@@ 7. Relation to other conjectures
In this section, we discuss the relationship between our conjecture and other conjectural statements, due to K. Saito and C. Hertling, which focus on the distribution of the spectral numbers.

### v3

@@ 3. Analytic torsion of $\mathcal{O}_{X}$
$\log\tau(\mathcal{X}_{t},\mathcal{O}_{\mathcal{X}_{t}})=(-1)^{n}\left(\frac{\mu}{(n+2)!}-\widetilde{p}_{g}\right)\log|t|^{2}+(-1)^{n}\beta\log\log|t|^{-1}+O(1),$
where $\beta$ is defined in (3.4). In particular:
  1. (1)
    If the weak form of the conjecture holds, then $\tau(\mathcal{X}_{t},\mathcal{O}_{\mathcal{X}_{t}})^{(-1)^{n}}$ converges to zero as $t$ goes to 0.
  2. (2)
    If the strong form of the conjecture holds, then $\tau(\mathcal{X}_{t},\mathcal{O}_{\mathcal{X}_{t}})^{(-1)^{n}}=O\left(|t|^{2m/(n+2)!-\varepsilon}\right)$ as $t\to 0$, for every $\varepsilon>0$. Here, $m$ is the number of singular points in $\mathcal{X}_{0}$.
  3. (3)
    If $\mathcal{X}_{0}$ has canonical singularities, then $\tau(\mathcal{X}_{t},\mathcal{O}_{\mathcal{X}_{t}})^{(-1)^{n}}=O(|t|^{2\mu/(n+2)!})$ as $t\to 0$.
Notice that, while the analytic torsion depends on the choice of metrics, the asymptotic behavior is metric independent. Also, for families of curves, the volume of the fibers with respect to a Kähler form on $\mathcal{X}$ is constant, and the $L^{2}$-norm on holomorphic differentials is independent of the choice of metric. Hence, the asymptotic behavior is then valid for any choice of Kähler metric on the total space $\mathcal{X}$. We remark that, for curves, $\tau(\mathcal{X}_{t},\mathcal{O}_{\mathcal{X}_{t}})^{-1}=\det\Delta^{0,1}_{\overline{\partial}}$ coincides with $\det\Delta^{0,0}_{\overline{\partial}}$, which in turn coincides with the determinant of the Laplace–Beltrami operator up to a topological constant.
@@
A variant of the above arises from the tautological family $\mathcal{H}$ of hypersurfaces of degree $d$ in $\mathbb{P}^{n+1}$ over the space of parameters $\mathbb{P}^{N}$, where $N=\binom{n+1}{d}-1$. Then $\mathcal{H}$ is a smooth space and inherits a Kähler metric from $\mathbb{P}^{n+1}$ and $\mathbb{P}^{N}$. Hence, the restriction along any curve $\Delta\to\mathbb{P}^{N}$ in which the total space is still smooth is an example of such a degeneration.
3.8.
Since any germ of isolated singularity admits a good compactification as in §2.8, we readily deduce, from Corollary 3.2, a reformulation of our conjecture in terms of analytic torsion.
Proposition 3.3.
For an isolated singularity $f=0$ of dimension $n\geq 1$, the following are equivalent:
  1. (1)
    The weak form, respectively strong form, of the conjecture holds for $f$.
  2. (2)
    For any good compactification $\mathcal{X}\to\Delta$ of $f$, and any Kähler form on $\mathcal{X}$ which is rational on smooth fibers, we have
    $\begin{split}&\tau(\mathcal{X}_{t},\mathcal{O}_{\mathcal{X}_{t}})^{(-1)^{n}}\to 0,\quad\text{as}\quad t\to 0,\\ &\text{respectively}\quad\tau(\mathcal{X}_{t},\mathcal{O}_{\mathcal{X}_{t}})^{(-1)^{n}}=O(|t|^{2/(n+2)!-\varepsilon}),\quad\text{for every }\quad\varepsilon>0,\quad\text{as}\quad t\to 0.\end{split}$
In the proposition above, one can replace any good compactification (resp. any Kähler form) by some good compactification (resp. some Kähler form). We also bring the reader’s attention to the fact that the positivity of $\beta$, established in Proposition 3.1, is fundamental to prove that (2) implies (1).
@@ 4. Quasi-homogeneous singularities
In this section, we address several cases of the strong form of the conjecture in the setting of quasi-homogeneous singularities. We also recall some useful notions about Newton polyhedra also utilized in later sections.
@@ 4.1. Newton polyhedra
where the sum runs over integers $k_{i}>0$ such that $\sum k_{i}w_{i}<1$. These are the interior lattice points of the Newton polyhedron of the polynomial $f(x_{0},\ldots,x_{n})$. Eventually, for convenience, we may include the lattice points with $\sum k_{i}w_{i}=1$, since these contribute zero to the sum. For the Milnor number, the formula (4.1) is given in [24, Theorem 1]. The expression (4.2) for $\widetilde{p}_{g}$ can be derived from (2.5) and the computation of the spectral polynomial of $f$ (cf. (2.6) for the definition), due to Steenbrink [37, Example 5.11]:
@@
To this end, it is enough to expand the polynomial into a power series of $T$ with rational exponents, and collect the terms whose exponents are strictly smaller than one. For this purpose, one can ignore the negative $T$ in the numerator, and expand $T^{w_{j}}/(1-T^{w_{j}})=T^{w_{j}}+T^{2w_{j}}+\ldots$. One finds that the exponents $\alpha^{\prime}$ strictly less than one are of the form $\alpha^{\prime}=\sum k_{i}w_{i}$, where the $k_{i}$ run over all the possible integers $k_{i}>0$ such that $\alpha^{\prime}<1$.
@@ 4.3. Homogeneous polynomials
It would be interesting to establish the strong form of the conjecture for quasi-homogeneous singularities, such as the Brieskorn–Pham singularities of the form
@@
corresponding to the choices of weights $w_{i}=1/a_{i}$. Here we treat the case when $f$ be a homogenenous polynomial in arbitrary dimension, so that all $w_{i}=\frac{1}{d}$. In this case (4.1) gives
@@
An inductive argument shows that the sum is a polynomial of degree $n+2$ in $d$. Moreover, the sum is empty for $d=0,\ldots,n+1,$ so the polynomial is of the form $C\cdot d(d-1)\cdots(d-(n+1))$ for some constant $C$. Since the sum is equal to $1$ for $d=n+2$ one sees that $C=\frac{1}{(n+2)!}$. ∎
@@ 4.4. Quasi-homogeneous singularities in dimension one
When $n=1$, the strong form of the conjecture is satisfied in the case of quasi-homogeneous singularities.
In preparation for the proof, we first notice that the invariants $\mu$ and $\widetilde{p}_{g}$ only depend on the analytical type of the germ of the singularity. By [31, Satz 1.3] we can assume that $w_{i}\leq\frac{1}{2}$. Then, by [48], the weights depend only on the topological type of the singularity, which is one of the following:
@@
The case when $a,b$ are relatively prime follows immediately from the computation of Mordell [25, Equation (4)]. The sum in (4.8) is
@@ 5. Irreducible curve singularities
See Némethi [28, Section 3], and in particular Theorem 3.1 therein, and M. Saito [35, Theorem 1.5 & Section 5]. For the spectral genus, we refer to §2.3 above for the expression in terms of M. Saito’s convention. ∎
@@ 6. Generic singularities with large Newton polyhedra
In this section, we consider isolated hypersurface singularities which are non-degenerate and convenient in the sense of Kouchnirenko [23]. [^2] In op. cit., which is written in French, convenient is called commode. These conditions are formulated in terms of the Newton polyhedron. The non-degeneracy condition is shown to be generic for the Zariski topology. Therefore, we may refer to such singularities as generic. Convenient means that the Newton polyhedron intersects all the coordinate hyperplanes. We refer to op. cit. for details.
@@
Consider now a convenient and generic isolated hypersurface singularity defined by $f\colon(\mathbb{C}^{n+1},0)\to(\mathbb{C},0)$. The formula of Kouchnirenko [23, Théorème I] states that
@@
The spectral genus can also be described explicitly in terms of the Newton polyhedron. Let $\Gamma_{c}$ be the compact Newton boundary. It is defined by a homogeneous, concave, piece-wise linear function $\phi$, which takes the value 1 on the faces of $\Gamma_{c}$. The lower Newton polyhedron can then be presented as $\Gamma_{-}=\{x\in\mathbb{R}_{+}^{n+1}\mid\phi(x)\leq 1\}$. By [34], in particular using the formulas on the first page, it follows that the spectral genus is given by
@@
In order to estimate an asymptotic form of the sum (6.2), we will apply a rough version of the Euler–Maclaurin formula for polyhedra. For the statement, we adopt the following convention for Lebesgue measures. Let $P$ be a convex lattice polyhedron in $\mathbb{R}^{d}$, and $Q$ a face of it. Let $\langle Q\rangle_{\mathbb{R}}\subseteq\mathbb{R}^{d}$ be the real vector space parallel to $Q$. Then we denote by $dx$ the Lebesgue measure on $\langle Q\rangle_{\mathbb{R}}$ which gives volume one to the fundamental domain of the lattice $\mathbb{Z}^{d}\cap\langle Q\rangle_{\mathbb{R}}$.
@@
This is an application of the local Euler–Maclaurin formula of Berline–Vergne [4, Section 5.4, Theorem 5]. The dominant term of the expansion is already given in loc. cit. They discuss the subdominant term in Section 5.5. Notice also that in [4] one supposes that $h$ is a smooth function on $\mathbb{R}^{d}$ with compact support. Since the sum and the first terms of the expansion depend only on the values of $h$ in $P$, which is compact, we can smoothly extend $h$ outside of $P$, still with compact support, for which the formula applies. ∎
@@
In general, the integrals appearing in the previous lemma can be evaluated by applying results of Brion [9, Section 3.2], see also [10, Corollary 6.1.10]. In our setting, it will be enough to have the following.
@@
which allows to conclude. ∎
@@
Let $f(x_{0},\ldots,x_{n})=0$ be a convenient and generic isolated singularity. Then, for $k$ sufficiently large, the isolated singularity $f(x_{0}^{k},\ldots,x_{n}^{k})=0$ satisfies the strong form of the conjecture.
@@
where $\phi$ is the piece-wise linear function determining the lower Newton polyhedron $\Gamma_{-}$. To estimate (6.5), we apply Lemma 6.1. For this, we first decompose $\Gamma_{-}$ into simplices with 0 as a vertex and the rest of the vertices in $\mathbb{Z}^{n+1}\cap\Gamma_{c}$. These define a simplicial complex, whose top dimensional simplices we denote by $\sigma_{i}$, and whose codimension 1 simplices we denote by $\tau_{j}$. On each simplex, the function $\phi$ is linear. A first application of the lemma gives
@@
For big enough $k$, the leading term, which is positive, dominates. This concludes the proof of the theorem. ∎
@@ 7. Relation to other conjectures
In this section, we discuss the relationship between our conjecture and other conjectural statements. We begin by showing that the weak conjecture can be derived from the Durfee-type conjecture. Then, we consider the relationship with the conjectures of K. Saito and C. Hertling, which focus on the distribution of the spectral numbers.
7.1. Durfee-type conjecture
Recall from the introduction the Durfee-type conjecture (1.1). In order to relate it to the weak conjecture, we begin by expressing the spectral genus in terms of the geometric genus of a suspension.
Let $f=0$ be an isolated hypersurface singularity of dimension $n\geq 1$, and recall that we denote by $T_{s}$ the semi-simple part of the monodromy acting on $\operatorname{Gr}_{F}^{n}H^{n}(\operatorname{Mil}_{f})$. Let $k\geq 1$ be an integer such that $T_{s}^{k}=1$. We define the $(k+1)$-suspension of $f$ as
$h(x_{0},\ldots,x_{n+1})=f(x_{0},\ldots,x_{n})+x_{n+1}^{k+1}.$
This function defines an isolated hypersurface singularity at 00, of dimension $n+1$.
Lemma 7.1.
With the notation as above, the spectral genus $\widetilde{p}_{g,f}$ of $f$ and the geometric genus $p_{g,h}$ of $h$ are related by
$p_{g,h}=k\cdot\widetilde{p}_{g,f}.$
Proof.
Let $\alpha_{i}^{\prime}$ be the spectral numbers of $f$, taken in $(0,n+1)$. Recall from §2.5 the Thom–Sebastiani property for the spectral numbers. The spectral numbers of $h$ are thus given by
$\beta_{i,j}=\alpha_{i}^{\prime}+\frac{j}{k+1},\quad\text{for}\ i=1,\ldots,\mu_{f}\ \text{and}\ j=1,\ldots,k.$
The geometric genus of $h$ equals #$\#\{\beta_{i,j}\leq 1\}$, by the very definition of the spectral numbers, cf. §2.3, and because $p_{g,h}=\dim H^{n+1}(\operatorname{Mil}_{h})$, cf. §2.2.
Define $\lambda_{i}=1-\alpha_{i}^{\prime}$. The condition $\beta_{i,j}\leq 1$ entails $\alpha_{i}^{\prime}<1$, and is actually equivalent to
(7.1) $j\leq k\lambda_{i}+\lambda_{i}.$
Because $T_{s}^{k}=1$, the quantities $k\lambda_{i}$ are integers. Moreover, $\lambda_{i}<1$. Therefore, given any $\alpha_{i}^{\prime}<1$, the equation (7.1) has exactly $k\lambda_{i}$ solutions. Recalling the expression (2.5) for $\widetilde{p}_{g,f}$, we find
$p_{g,h}=\#\{\beta_{i,j}\leq 1\}=\sum_{\alpha_{i}^{\prime}<1}k(1-\alpha_{i}^{\prime})=k\cdot\widetilde{p}_{g,f}.$
This concludes the proof. ∎
Proposition 7.2.
With the notation as above, the weak conjecture for $f$ is equivalent to the Durfee-type conjecture for $h$.
Proof.
The Milnor number of $h$ is $\mu_{h}=k\cdot\mu_{f}$. Hence, by the previous lemma, the Durfee-type conjecture for $h$ amounts to
$k\cdot\widetilde{p}_{g,f}<\frac{k\cdot\mu_{f}}{(n+2)!},$
which is equivalent to the weak conjecture for $f$. ∎
7.2.
We now apply the previous proposition to address Theorem B in the introduction.
Proof of Theorem B. The suspension of a quasi-homogeneous singularity is again a quasi-homogeneous singularity. For these, the Durfee-type conjecture is a result of Yau–Zhang [45, Theorem 1.2]. We conclude by Proposition 7.2.
For the case of isolated plane curve singularities, Némethi studies in [29] the Durfee conjecture for suspensions of the form $h(x,y,z)=f(x,y)+z^{N}$. He requires that $N\geq 2$ is coprime with the multiplicities of the exceptional divisors of the minimal embedded resolution of singularities of $f=0$. We can apply his results with $N=k+1$, where $k$ is any multiple of the least common multiple of the multiplicities of the exceptional divisors. Then $T_{s}^{k}=1$ by [1, Théorème 3], so that we are in the setting of Proposition 7.2. In [29, Theorem 5.1], it is proven that
$p_{g,h}\leq\frac{\mu_{h}}{6}.$
We will check that, apart possibly from some exceptional cases, this is actually a strict inequality. The exceptional cases will be dealt with separately.
Let $\sigma$ be the signature of the intersection form on $H^{2}(\operatorname{Mil}_{h},\mathbb{Z})$. Némethi observes that, for $N$ chosen as above, we have $\sigma+\mu_{h}=4p_{g,h}$. We thus need to show that $\sigma<-\mu_{h}/3$. If $f$ is not equivalent to a singularity of the form $A_{n}$, $A_{2n,2m}$, $D_{2n+3}$ or $E_{6}$, then, recalling that $\mu_{h}=(N-1)\mu_{f}$, the equation in the statement of [29, Theorem 5.1 (d)] gives
$\sigma\leq-\frac{N^{2}-1}{3N}\mu_{f}=-\frac{N+1}{3N}\mu_{h}<-\frac{\mu_{h}}{3}.$
Indeed, under the assumption on $f$, the quantity $\epsilon_{f}$ in that statement is 0 by definition.
If $f$ is equivalent to $A_{n}$, $D_{2n+3}$ or $E_{6}$, then the strong form of the conjecture holds by Theorem 4.3, since these are all quasi-homogeneous singularities.
For a singularity of type $A_{2n,2m}$, Némethi’s treatment shows that $\sigma<-\mu_{h}/3$ if $n+m\geq 5$, but it fails to provide a strict inequality otherwise. Instead, we deliver a simpler argument with a better outcome: the strong conjecture holds as well. Recall first that this singularity has equation
$f(x,y)=(x^{2}+y^{2n+1})(y^{2}+x^{2m+1}).$
The lower Newton polygon has vertices at the points $(0,0)$, $A=(0,2n+1)$, $B=(2,2)$ and $C=(2m+1,0)$. It is thus convenient, and the compact Newton boundary has two edges, namely the segments $AB$ and $BC$. It is straightforward to see that the principal parts of $f$ corresponding to $AB$ and $BC$ have non-vanishing gradient over (×$(\mathbb{C}^{\times})^{2}$, so that $f$ is non-degenerate in the sense of Kouchnirenko. Consequently, the discussion in §6.1 applies, and the spectral genus is given by the formula (6.2). The evaluation of the latter is elementary, and we omit the details. The result is
$\widetilde{p}_{g,f}=\frac{(n+2)^{2}}{2(2n+3)}+\frac{(m+2)^{2}}{2(2m+3)}-\frac{1}{2}.$
As for the Milnor number, Kouchnirenko’s formula (6.1) yields $\mu_{f}=2n+2m+7$. With these expressions at hand, we find
$\frac{\mu_{f}}{6}-\widetilde{p}_{g,f}\geq\frac{8}{15}>\frac{1}{6}.$
This concludes the proof.